# Effect of flow-aligned external magnetic fields on mushroom instability

### v1

@@ IV Conclusion
In this paper, we studied the effect of a flow-aligned magnetic field on MI, with both theoretical analyses and PIC simulations. In the theoretical part, we consider a generic shear flow under a uniform magnetic field, and analyze the system with two-fluid equations in the limit of a cold and collisionless plasma. In this way, we derive the eigenequations for the linear growth rates of MI. With a extension on the perturbation wavevector, our eigenequation can be generalized to describe not only MI (with k perpendicular to the shear), but also ESKHI (k parallel to the flow) and a more general coupled instability with an arbitrary wavevector k in the plane of the shear flow.
@@
Numerical solutions of the MI dispersion relation suggest that the external flow-aligned magnetic field always has a suppressing effect on MI. However, compared with ESKHI, MI is more robust to the external magnetic field, and in our framework there seems no stabilizing threshold magnetic field for MI; In other words, MI remains unstable under very strong magnetic fields. In consequence, we expect MI to dominate the dynamics of the shear flow not only in relativistic scenarios, but also in strongly magnetized sub-relativistic scenarios. We also investigate a few cases of the coupled instability with k over the whole plane. The result reveals that in shear flows with $V_{0}\sim c/3$, the coupling can lead to a dynamics different from a pure ESKHI or MI, and the external magnetic field can be destabilizing for some wavevectors. On contrary, in relativistic scenarios the instability growth rate is still peaked at $k_{y}=0$, which corresponds to a pure MI.
@@
In the simulation part, we investigate MI in a planar and collisionless shear flow with PIC simulations. First, we conduct single-mode simulations in cold plasmas ($v_{th}=0$) under different magnitudes of external magnetic fields. Results of the simulations reach a good agreement with our theoretical results, where growth rates extracted from the simulations are very close to our numerical solution of the MI dispersion relation. In unmagnetized runs, we observe the formation of the mushrooms and a dramatical deformation of the flow bulk in the saturation stage. On contrary, in magnetized runs the mushrooms are redirected to gyrate in the plane, restricting the deformation of the flow bulk. Especially, a quasi-steady density gap is formed in the magnetized runs, where electrons and protons are pushed away from the shear interface, indicating the potential of magnetized MI in jet spine collimation.
@@
Simulations with finite temperatures are also carried out. In thermal simulations, MI growth spontaneously from thermal fluctuations, and merge into each other as they enter the nonlinear stage. As temperature increases, a diffusion-induced DC magnetic field occurs on the shear interface, which shares a similar morphology with the MI saturation field. This DC field grows faster than MI and dominates the early stage of the shear flow. In shear flows with higher temperatures, the DC field can achieve a high magnitude, hence skips or suppresses the linear growth of MI.
@@
Observations have revealed that the magnetic field structure can be actually helical in astrophysical jets (Hovatta et al., 2012). How this magnetic field interacts, or originates from some mechanism associated with MI, remains a critical open question. Since MI has only been discovered recently (Alves et al., 2015) and received limited attention in mainstream jet research, further work is needed to bridge this plasma instability from theory to large-scale and realistic jet dynamics.
Acknowledgements.
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDA25010100 and XDA25050500), National Natural Science Foundation of China (Grant No. 12075204) and Shanghai Municipal Science and Technology Key Project (Grant No. 22JC1401500). Dong Wu thanks the sponsorship from Yangyang Development Fund.
Appendix A 3D shear instability analyses
In this appendix we show the linearization process of Eqs. 1 $\sim$ 5 in detail. The physical settings of the shear flow are the same as in Sec. II.1, but the perturbations are generalized to be in the form of $f(x,y,z,t)=\tilde{f}(x)e^{i(k_{y}y+k_{z}z-\omega t)}$. The linearized form of Eqs. 1 and 2 is:
$\tilde{n}_{e}=\frac{-i}{\Omega}\left[\frac{\partial}{\partial x}(\tilde{v}_{x}n_{e0})+in_{e0}(k_{y}\tilde{v}_{y}+k_{z}\tilde{v}_{z})\right],$ (13)
$\tilde{v}_{x}=\frac{e}{\gamma_{0}m_{e}}\frac{-i}{\Omega}(\tilde{E}_{x}+v_{0}\tilde{B}_{z}-\tilde{v}_{z}B_{0}),$ (14)
$\tilde{v}_{y}=\frac{e}{\gamma_{0}^{3}m_{e}}\frac{-i}{\Omega}\left[\tilde{E}_{y}+\frac{m_{e}}{e}\tilde{v}_{x}\frac{\partial}{\partial x}(\gamma_{0}v_{0})\right],$ (15)
$\tilde{v}_{z}=\frac{e}{\gamma_{0}m_{e}}\frac{-i}{\Omega}(\tilde{E}_{z}-v_{0}\tilde{B}_{x}+\tilde{v}_{x}B_{0}),$ (16)
where $\Omega\equiv\omega-k_{y}v_{0}$. We can decouple $\tilde{v}_{x}$ and $\tilde{v}_{z}$ from Eqs. 14 and 16:
$\tilde{v}_{x}=-\frac{e}{\gamma_{0}m_{e}}\frac{\omega_{c}(-\tilde{E}_{z}+\tilde{B}_{x}v_{0})+i\Omega(\tilde{E}_{x}+\tilde{B}_{z}v_{0})}{\Omega^{2}-\omega_{c}^{2}},$ (17)
$\tilde{v}_{z}=-\frac{e}{\gamma_{0}m_{e}}\frac{\omega_{c}(\tilde{E}_{x}+\tilde{B}_{z}v_{0})+i\Omega(\tilde{E}_{z}-\tilde{B}_{x}v_{0})}{\Omega^{2}-\omega_{c}^{2}}.$ (18)
@@ Acknowledgements.
$\tilde{J}_{x}=\frac{e^{2}n_{e0}}{\gamma_{0}m_{e}}\frac{\omega_{c}(-\tilde{E}_{z}+\tilde{B}_{x}v_{0})+i\Omega(\tilde{E}_{x}+\tilde{B}_{z}v_{0})}{\Omega^{2}-\omega_{c}^{2}},$ (19)
$\tilde{J}_{y}=\frac{e^{2}n_{e0}}{\gamma_{0}^{3}m_{e}}\frac{i\omega}{\Omega^{2}}\tilde{E}_{y}+ie\frac{\partial}{\partial x}\left(\frac{v_{0}\tilde{v}_{x}n_{e0}}{\Omega}\right)-\frac{ev_{0}n_{e0}}{\Omega}k_{z}\tilde{v}_{z},$ (20)
$\tilde{J}_{z}=\frac{e^{2}n_{e0}}{\gamma_{0}m_{e}}\frac{\omega_{c}(\tilde{E}_{x}+\tilde{B}_{z}v_{0})+i\Omega(\tilde{E}_{z}-\tilde{B}_{x}v_{0})}{\Omega^{2}-\omega_{c}^{2}},$ (21)
@@
$\tilde{B}_{x}=\frac{1}{\omega}(k_{y}\tilde{E}_{z}-k_{z}\tilde{E}_{y}),$ (22)
$\tilde{B}_{z}=-\frac{i}{\omega}\frac{\partial\tilde{E}_{y}}{\partial x}-\frac{k_{y}}{\omega}\tilde{E}_{x},$ (23)
Replace B in Eq. 4 with Eqs. 22 and 23, we can derive
$ik_{y}\frac{\partial\tilde{E}_{y}}{\partial x}+ik_{z}\frac{\partial\tilde{E}_{z}}{\partial x}=i\omega\mu_{0}\tilde{J}_{x}+\left(\frac{\omega^{2}}{c^{2}}-k_{y}^{2}-k_{z}^{2}\right)\tilde{E}_{x},$ (24)
$ik_{y}\frac{\partial\tilde{E}_{x}}{\partial x}-\frac{\partial^{2}\tilde{E}_{y}}{\partial x^{2}}-k_{y}k_{z}\tilde{E}_{z}=i\omega\mu_{0}\tilde{J}_{y}+\left(\frac{\omega^{2}}{c^{2}}-k_{z}^{2}\right)\tilde{E}_{y},$ (25)
$ik_{z}\frac{\partial\tilde{E}_{x}}{\partial x}-\frac{\partial^{2}\tilde{E}_{z}}{\partial x^{2}}-k_{y}k_{z}\tilde{E}_{y}=i\omega\mu_{0}\tilde{J}_{z}+\left(\frac{\omega^{2}}{c^{2}}-k_{y}^{2}\right)\tilde{E}_{z}.$ (26)
Since $\tilde{J}_{x}$ only contains first order terms of $\tilde{E}_{x}$, from Eq. 24 we can solve $\tilde{E}_{x}$ in terms of only two perturbed variables $\tilde{E}_{y}$ and $\tilde{E}_{z}$:
$\tilde{E}_{x}=$ $\frac{i\left[k_{y}c^{2}\left(\Omega^{2}-\omega_{c}^{2}\right)-\gamma_{0}^{2}\omega_{p}^{2}\Omega v_{0}\right]}{D(\omega,k)}\frac{\partial\tilde{E}_{y}}{\partial x}+\frac{ik_{z}c^{2}\left(\Omega^{2}-\omega_{c}^{2}\right)}{D(\omega,k)}\frac{\partial\tilde{E}_{z}}{\partial x}+\frac{i\gamma_{0}^{2}\omega_{p}^{2}\omega_{c}k_{z}v_{0}}{D(\omega,k)}\tilde{E}_{y}+\frac{i\gamma_{0}^{2}\omega_{p}^{2}\omega_{c}\Omega}{D(\omega,k)}\tilde{E}_{z},$ (27)
where $D(\omega,k)=\Omega^{2}\left[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2}\right]-\omega_{c}^{2}[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}]$ is the generalization of $D(\omega,k_{z})$ defined in Sec. II.2. Now all perturbed variables in the system can be expressed as functions of $\tilde{E}_{y},\tilde{E}_{z}$ and their partial derivatives with respect to $x$. Substitute them into Eqs. 25 and 26, we finally derive the two coupled eigenequations of the system:
@@
$+\left[\frac{\omega_{c}\gamma_{0}^{2}\omega_{p}^{2}\left(\omega^{2}v_{0}/c^{2}-\omega k_{y}-k_{z}^{2}v_{0}\right)}{D(\omega,k)}\tilde{E}_{z}\right]^{\prime}+\frac{k_{z}^{2}v_{0}\gamma_{0}^{2}\omega_{c}\omega_{p}^{2}}{D(\omega,k)}\tilde{E}_{z}^{\prime}+\left[\frac{\Omega k_{z}v_{0}\gamma_{0}^{2}\omega_{p}^{2}}{c^{2}(\Omega^{2}-\omega_{c}^{2})}-k_{y}k_{z}\right]\tilde{E}_{z}=0,$ (28)
@@
$-\frac{\omega_{c}\gamma_{0}^{2}\omega_{p}^{2}\left(k_{y}\omega+k_{z}^{2}v_{0}-\omega^{2}v_{0}/c^{2}\right)}{D(\omega,k)}\tilde{E}_{y}^{\prime}+\left[\frac{k_{z}v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}(c^{2}k^{2}-\omega^{2}+\gamma_{0}^{2}\omega_{p}^{2})}{c^{2}D(\omega,k)}+k_{y}k_{z}\right]\tilde{E}_{y}=0,$ (29)
where $F(\omega,k)\equiv\Omega^{2}[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2}]^{2}-\omega_{c}^{2}[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}]^{2}$. Still, the connecting boundary conditions are derived by integrating Eqs. 28 and 29 across the shear interface (e.g. $x=0$):
@@
$+\left[\frac{k_{z}(v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}-k_{y}c^{2}(\Omega^{2}-\omega_{c}^{2}))}{D(\omega,k)}\tilde{E}_{z}^{\prime}\right]\bigg|_{0_{-}}^{0_{+}}+\left[\frac{\omega_{c}\gamma_{0}^{2}\omega_{p}^{2}\left(\omega^{2}v_{0}/c^{2}-\omega k_{y}-k_{z}^{2}v_{0}\right)}{D(\omega,k)}\tilde{E}_{z}\right]\bigg|_{0_{-}}^{0_{+}}=0,$ (30)
@@
$+\left[\frac{k_{z}[c^{2}k_{y}(\Omega^{2}-\omega_{c}^{2})-v_{0}\gamma_{0}^{2}\Omega\omega_{p}^{2}]}{D(\omega,k)}\tilde{E}_{y}^{\prime}\right]\bigg|_{0_{-}}^{0_{+}}+\left[\frac{k_{z}^{2}v_{0}\gamma_{0}^{2}\omega_{c}\omega_{p}^{2}}{D(\omega,k)}\tilde{E}_{y}\right]\bigg|_{0_{-}}^{0_{+}}=0,$ (31)
@@
Instead, in the unmagnetized case with $\omega_{c}=0$, Eqs. 28 and 29 become
@@
$+\left[\frac{k_{z}(v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}-k_{y}c^{2}\Omega^{2})}{\Omega^{2}\left[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2}\right]}\tilde{E}_{z}^{\prime}\right]^{\prime}+\left[\frac{k_{z}(v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}-k_{y}c^{2}\Omega^{2})}{c^{2}\Omega^{2}}\right]\tilde{E}_{z}=0,$ (32)
@@
$+\left[\frac{k_{z}(c^{2}k_{y}\Omega^{2}-v_{0}\gamma_{0}^{2}\Omega\omega_{p}^{2})}{\Omega^{2}(\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2})}\tilde{E}_{y}^{\prime}\right]^{\prime}+\left[\frac{k_{z}(c^{2}k_{y}\Omega^{2}-v_{0}\gamma_{0}^{2}\Omega\omega_{p}^{2})}{c^{2}\Omega^{2}}\right]\tilde{E}_{y}=0,$ (33)
where one can easily find that the solution to $\tilde{E}_{y}$ and $\tilde{E}_{z}$ is the single mode $C_{i}\exp\left[-|x|\sqrt{(k_{y}^{2}+k_{z}^{2})+(\gamma_{0}^{2}\omega_{p}^{2}-\omega^{2})/c^{2}}\right]$. Substitute the solution into Eqs. 30 and 31, and one can derive the dispersion relation of the unmagnetized shear flow. The special case of unmagnetized ESKHI (with $k_{z}=0$) and MI (with $k_{y}=0$) can be solved analytically.

### v2

@@ 4 Conclusion
In this paper, we have investigated the effect of a flow-aligned magnetic field on MI, with both theoretical analyses and PIC simulations. In the theoretical part, we consider a generic shear flow under a uniform magnetic field, and analyze the system with two-fluid equations in the limit of a cold and collisionless plasma. In this way, we derive the eigenequations for the linear growth rates of MI. With a extension on the perturbation wavevector, our eigenequation can be generalized to describe not only MI (with k perpendicular to the shear), but also ESKHI (k parallel to the flow) and a more general coupled instability with an arbitrary wavevector k in the plane of the shear flow.
@@
Numerical solutions of the MI dispersion relation suggest that the external flow-aligned magnetic field always has a suppressing effect on MI. However, compared with ESKHI, MI is more robust against the external magnetic field, and within our framework there appears to be no stabilizing threshold magnetic field for MI; In other words, MI remains unstable under very strong magnetic fields. In consequence, we expect MI to dominate the dynamics of the shear flow not only in relativistic scenarios, but also in strongly magnetized sub-relativistic scenarios. We also investigate a few cases of the coupled instability with k over the whole plane. The result reveals that in shear flows with $V_{0}\sim c/3$, the coupling can lead to a dynamics different from a pure ESKHI or MI, and the external magnetic field can be destabilizing for some wavevectors. In contrast, in relativistic scenarios the instability growth rate still peaks at $k_{y}=0$, which corresponds to a pure MI.
@@
In the simulation part, we investigate MI in a planar and collisionless shear flow with PIC simulations. First, we conduct single-mode simulations in cold plasmas ($v_{th}=0$) under different magnitudes of external magnetic fields. Results of the simulations reach a good agreement with our theoretical results, where growth rates extracted from the simulations are very close to our numerical solution of the MI dispersion relation. In unmagnetized runs, we observe the formation of the mushrooms and a dramatical deformation of the flow bulk in the saturation stage. In contrast, in magnetized runs the mushrooms are redirected to gyrate in the plane, restricting the deformation of the flow bulk. Especially, a quasi-steady density gap is formed in the magnetized runs, where electrons and protons are pushed away from the shear interface, indicating the potential of magnetized MI in jet spine collimation.
@@
Simulations with finite temperatures are also carried out. In thermal simulations, MI growth spontaneously from thermal fluctuations, and the “mushrooms” merge into each other as they enter the nonlinear stage. As temperature increases, a diffusion-induced DC magnetic field occurs on the shear interface, which shares a similar morphology with the MI saturation field. This DC field grows faster than MI and dominates the early stage of the shear flow. In shear flows with higher temperatures, the DC field can reach a high magnitude, thereby skipping or suppressing the linear growth phase of MI.
@@
Observations have revealed that the magnetic field structure can be actually helical in astrophysical jets (Hovatta et al., 2012). How this magnetic field interacts, or originates from some mechanism associated with MI, remains a critical open question. Moreover, how the cylindrical geometry of jets affects MI remains theoretically unclear. Since MI has only been discovered recently (Alves et al., 2015) and received limited attention in mainstream jet research, further work is needed to bridge this plasma instability from simplified theory and simulations to large-scale and realistic jet dynamics.
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDA25010100 and XDA25050500), Science Challenge Project (No. TZ2025012), National Natural Science Foundation of China (Grant No. 12075204) and Shanghai Municipal Science and Technology Key Project (Grant No. 22JC1401500). Dong Wu thanks the sponsorship from Yangyang Development Fund.
Appendix A generalized shear instability analyses
In this appendix we show the linearization process of Eqs. 1 $\sim$ 5 in detail. The physical settings of the shear flow are the same as in Sec. 2.1, but the perturbations are generalized to be in the form of $f(x,y,z,t)=\tilde{f}(x)e^{i(k_{y}y+k_{z}z-\omega t)}$. The linearized form of Eqs. 1 and 2 is:
$\tilde{n}_{e}=\frac{-i}{\Omega}\left[\frac{\partial}{\partial x}(\tilde{v}_{x}n_{e0})+in_{e0}(k_{y}\tilde{v}_{y}+k_{z}\tilde{v}_{z})\right],$ (A1)
$\tilde{v}_{x}=\frac{e}{\gamma_{0}m_{e}}\frac{-i}{\Omega}(\tilde{E}_{x}+v_{0}\tilde{B}_{z}-\tilde{v}_{z}B_{0}),$ (A2)
$\tilde{v}_{y}=\frac{e}{\gamma_{0}^{3}m_{e}}\frac{-i}{\Omega}\left[\tilde{E}_{y}+\frac{m_{e}}{e}\tilde{v}_{x}\frac{\partial}{\partial x}(\gamma_{0}v_{0})\right],$ (A3)
$\tilde{v}_{z}=\frac{e}{\gamma_{0}m_{e}}\frac{-i}{\Omega}(\tilde{E}_{z}-v_{0}\tilde{B}_{x}+\tilde{v}_{x}B_{0}),$ (A4)
where $\Omega\equiv\omega-k_{y}v_{0}$. We can decouple $\tilde{v}_{x}$ and $\tilde{v}_{z}$ from Eqs. A2 and A4:
$\tilde{v}_{x}=-\frac{e}{\gamma_{0}m_{e}}\frac{\omega_{c}(-\tilde{E}_{z}+\tilde{B}_{x}v_{0})+i\Omega(\tilde{E}_{x}+\tilde{B}_{z}v_{0})}{\Omega^{2}-\omega_{c}^{2}},$ (A5)
$\tilde{v}_{z}=-\frac{e}{\gamma_{0}m_{e}}\frac{\omega_{c}(\tilde{E}_{x}+\tilde{B}_{z}v_{0})+i\Omega(\tilde{E}_{z}-\tilde{B}_{x}v_{0})}{\Omega^{2}-\omega_{c}^{2}}.$ (A6)
@@
$\tilde{J}_{x}=\frac{e^{2}n_{e0}}{\gamma_{0}m_{e}}\frac{\omega_{c}(-\tilde{E}_{z}+\tilde{B}_{x}v_{0})+i\Omega(\tilde{E}_{x}+\tilde{B}_{z}v_{0})}{\Omega^{2}-\omega_{c}^{2}},$ (A7)
$\tilde{J}_{y}=\frac{e^{2}n_{e0}}{\gamma_{0}^{3}m_{e}}\frac{i\omega}{\Omega^{2}}\tilde{E}_{y}+ie\frac{\partial}{\partial x}\left(\frac{v_{0}\tilde{v}_{x}n_{e0}}{\Omega}\right)-\frac{ev_{0}n_{e0}}{\Omega}k_{z}\tilde{v}_{z},$ (A8)
$\tilde{J}_{z}=\frac{e^{2}n_{e0}}{\gamma_{0}m_{e}}\frac{\omega_{c}(\tilde{E}_{x}+\tilde{B}_{z}v_{0})+i\Omega(\tilde{E}_{z}-\tilde{B}_{x}v_{0})}{\Omega^{2}-\omega_{c}^{2}},$ (A9)
@@
$\tilde{B}_{x}=\frac{1}{\omega}(k_{y}\tilde{E}_{z}-k_{z}\tilde{E}_{y}),$ (A10)
$\tilde{B}_{z}=-\frac{i}{\omega}\frac{\partial\tilde{E}_{y}}{\partial x}-\frac{k_{y}}{\omega}\tilde{E}_{x},$ (A11)
Replace B in Eq. 4 with Eqs. A10 and A11, we can derive
$ik_{y}\frac{\partial\tilde{E}_{y}}{\partial x}+ik_{z}\frac{\partial\tilde{E}_{z}}{\partial x}=i\omega\mu_{0}\tilde{J}_{x}+\left(\frac{\omega^{2}}{c^{2}}-k_{y}^{2}-k_{z}^{2}\right)\tilde{E}_{x},$ (A12)
$ik_{y}\frac{\partial\tilde{E}_{x}}{\partial x}-\frac{\partial^{2}\tilde{E}_{y}}{\partial x^{2}}-k_{y}k_{z}\tilde{E}_{z}=i\omega\mu_{0}\tilde{J}_{y}+\left(\frac{\omega^{2}}{c^{2}}-k_{z}^{2}\right)\tilde{E}_{y},$ (A13)
$ik_{z}\frac{\partial\tilde{E}_{x}}{\partial x}-\frac{\partial^{2}\tilde{E}_{z}}{\partial x^{2}}-k_{y}k_{z}\tilde{E}_{y}=i\omega\mu_{0}\tilde{J}_{z}+\left(\frac{\omega^{2}}{c^{2}}-k_{y}^{2}\right)\tilde{E}_{z}.$ (A14)
Since $\tilde{J}_{x}$ only contains first order terms of $\tilde{E}_{x}$, from Eq. A12 we can solve $\tilde{E}_{x}$ in terms of only two perturbed variables $\tilde{E}_{y}$ and $\tilde{E}_{z}$:
$\tilde{E}_{x}=$ $\frac{i\left[k_{y}c^{2}\left(\Omega^{2}-\omega_{c}^{2}\right)-\gamma_{0}^{2}\omega_{p}^{2}\Omega v_{0}\right]}{D(\omega,k)}\frac{\partial\tilde{E}_{y}}{\partial x}+\frac{ik_{z}c^{2}\left(\Omega^{2}-\omega_{c}^{2}\right)}{D(\omega,k)}\frac{\partial\tilde{E}_{z}}{\partial x}+\frac{i\gamma_{0}^{2}\omega_{p}^{2}\omega_{c}k_{z}v_{0}}{D(\omega,k)}\tilde{E}_{y}+\frac{i\gamma_{0}^{2}\omega_{p}^{2}\omega_{c}\Omega}{D(\omega,k)}\tilde{E}_{z},$ (A15)
where $D(\omega,k)=\Omega^{2}\left[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2}\right]-\omega_{c}^{2}[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}]$ is the generalization of $D(\omega,k_{z})$ defined in Sec. 2.2. Now all perturbed variables in the system can be expressed as functions of $\tilde{E}_{y},\tilde{E}_{z}$ and their partial derivatives with respect to $x$. Substitute them into Eqs. A13 and A14, we finally derive the two coupled eigenequations of the system:
@@
$+\left[\frac{\omega_{c}\gamma_{0}^{2}\omega_{p}^{2}\left(\omega^{2}v_{0}/c^{2}-\omega k_{y}-k_{z}^{2}v_{0}\right)}{D(\omega,k)}\tilde{E}_{z}\right]^{\prime}+\frac{k_{z}^{2}v_{0}\gamma_{0}^{2}\omega_{c}\omega_{p}^{2}}{D(\omega,k)}\tilde{E}_{z}^{\prime}+\left[\frac{\Omega k_{z}v_{0}\gamma_{0}^{2}\omega_{p}^{2}}{c^{2}(\Omega^{2}-\omega_{c}^{2})}-k_{y}k_{z}\right]\tilde{E}_{z}=0,$ (A16)
@@
$-\frac{\omega_{c}\gamma_{0}^{2}\omega_{p}^{2}\left(k_{y}\omega+k_{z}^{2}v_{0}-\omega^{2}v_{0}/c^{2}\right)}{D(\omega,k)}\tilde{E}_{y}^{\prime}+\left[\frac{k_{z}v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}(c^{2}k^{2}-\omega^{2}+\gamma_{0}^{2}\omega_{p}^{2})}{c^{2}D(\omega,k)}+k_{y}k_{z}\right]\tilde{E}_{y}=0,$ (A17)
where $F(\omega,k)\equiv\Omega^{2}[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2}]^{2}-\omega_{c}^{2}[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}]^{2}$. Still, the connecting boundary conditions are derived by integrating Eqs. A16 and A17 across the shear interface (e.g. $x=0$):
@@
$+\left[\frac{k_{z}(v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}-k_{y}c^{2}(\Omega^{2}-\omega_{c}^{2}))}{D(\omega,k)}\tilde{E}_{z}^{\prime}\right]\bigg|_{0_{-}}^{0_{+}}+\left[\frac{\omega_{c}\gamma_{0}^{2}\omega_{p}^{2}\left(\omega^{2}v_{0}/c^{2}-\omega k_{y}-k_{z}^{2}v_{0}\right)}{D(\omega,k)}\tilde{E}_{z}\right]\bigg|_{0_{-}}^{0_{+}}=0,$ (A18)
@@
$+\left[\frac{k_{z}[c^{2}k_{y}(\Omega^{2}-\omega_{c}^{2})-v_{0}\gamma_{0}^{2}\Omega\omega_{p}^{2}]}{D(\omega,k)}\tilde{E}_{y}^{\prime}\right]\bigg|_{0_{-}}^{0_{+}}+\left[\frac{k_{z}^{2}v_{0}\gamma_{0}^{2}\omega_{c}\omega_{p}^{2}}{D(\omega,k)}\tilde{E}_{y}\right]\bigg|_{0_{-}}^{0_{+}}=0,$ (A19)
@@
Instead, in the unmagnetized case with $\omega_{c}=0$, Eqs. A16 and A17 become
@@
$+\left[\frac{k_{z}(v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}-k_{y}c^{2}\Omega^{2})}{\Omega^{2}\left[\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2}\right]}\tilde{E}_{z}^{\prime}\right]^{\prime}+\left[\frac{k_{z}(v_{0}\Omega\gamma_{0}^{2}\omega_{p}^{2}-k_{y}c^{2}\Omega^{2})}{c^{2}\Omega^{2}}\right]\tilde{E}_{z}=0,$ (A20)
@@
$+\left[\frac{k_{z}(c^{2}k_{y}\Omega^{2}-v_{0}\gamma_{0}^{2}\Omega\omega_{p}^{2})}{\Omega^{2}(\omega^{2}-(k_{y}^{2}+k_{z}^{2})c^{2}-\gamma_{0}^{2}\omega_{p}^{2})}\tilde{E}_{y}^{\prime}\right]^{\prime}+\left[\frac{k_{z}(c^{2}k_{y}\Omega^{2}-v_{0}\gamma_{0}^{2}\Omega\omega_{p}^{2})}{c^{2}\Omega^{2}}\right]\tilde{E}_{y}=0,$ (A21)
where one can easily find that the solution to $\tilde{E}_{y}$ and $\tilde{E}_{z}$ is the single mode $C_{i}\exp\left[-|x|\sqrt{(k_{y}^{2}+k_{z}^{2})+(\gamma_{0}^{2}\omega_{p}^{2}-\omega^{2})/c^{2}}\right]$. Substitute the solution into Eqs. A18 and A19, and one can derive the dispersion relation of the unmagnetized shear flow. The special case of unmagnetized ESKHI (with $k_{z}=0$) and MI (with $k_{y}=0$) can be solved analytically.